OpenStudy (anonymous):

Find the set notation and interval notation of -3<=x<17

OpenStudy (anonymous):

Interval notation: [-3,17) Set notation: what's that

OpenStudy (anonymous):

idk?

OpenStudy (anonymous):

 That is set notation

OpenStudy (anonymous):

OH! {x \in R| -3 <=x <17}

OpenStudy (anonymous):

Can you draw it because im confused

OpenStudy (anonymous):

@mathstudent55

OpenStudy (solomonzelman):

\(\Large \left\{\left. x \in {\rm R} \right| -3\le x<17\right\}\) Translating what flyingpie said in latex. \(\Large \left\{\left. \color{blue}{x \in {\rm R} }\right| \color{red}{-3\le x<17}\right\}\) The notation in blue tells us that "x belongs to a set of all Real numbers". In other words x can be any real number. The notation in red tells us, that x can be -3 or larger, but is less than 17.

OpenStudy (anonymous):

Thanks :)! Can I ask one more?

OpenStudy (solomonzelman):

Square bracket indicates that we include the value Round bracket indicates that the value is not included So for example if I had: \(-1>x>5\) then my interval would be \((-1,5)\) if i had: \(-1<x\le5\) then my interval would be \((-1,5]\) if i had: \(-1\le x<5\) then my interval would be \([-1,5)\) if i had: \(-1\le x\le5\) then my interval would be \([-1,5]\)

OpenStudy (solomonzelman):

I have to go right now, sorry-:(

OpenStudy (anonymous):

Aw :( Okay